### 07 Jan 2024 v1

# On Some Quantum Correction to the Coulomb Potential in Generalized Uncertainty Principle Approach

Abstract
Taking into account the importance of the unified theory of quantum mechanics and gravity, and the existence of a minimum length of the order of the Planck scale, we consider a modified Schrödinger equation resulting from a generalised uncertainty principle, which finds applications from the realm of quantum information to large-scale physics, with a quantum mechanically corrected gravitational interaction proposed very recently. As the resulting equation cannot be solved by common exact approaches, including Heun or Lie algebraic ones, we propose a Bethe-Ansatz approach, which will be applied and whose results we discuss.
minimal length, generalized uncertainty principle, Planck scale, Schrödinger equation, quantum correction, Coulomb potential
PACS: 04.60.-m, 03.65.Ge.
I Introduction
There have been several theories to unifyingly describing quantum mechanics and gravity, including quantum gravity, black hole physics, double special relativity, and string theory, among others (see Konishi ; Maggiore ; Cortes ; Hossenfelder ; Scardigli and references quoted therein), with each approach having its own strengths and failures. However, the common point in such theories is that they all predict the existence of the so-called minimum length (ML) which is on the Planck scale $\ell_{p}=\sqrt{\hbar G/c^{3}}$. This ML is equivalent to a generalization of the uncertainty principle (GUP) that affects the entire physical system.
From the mathematical point of view, in a GUP formalism the equations become more complicated in structure and finding analytical solutions becomes a really challenging task. In particular, what in these circumstances we could call the Schrödinger equation appears in higher order forms: four, six or more, depending on the choice of the GUP and the operators involved. Obviously, the latter is not a very well studied category in mathematical physics, where in general the differential equations that appear, whether relativistic or non-relativistic, are of the first or second order.
As a consequence, considerable efforts have been made to investigate various aspects of GUP formalism. For example, Ali et al. Ali proposed a GUP compatible with string theory, black hole physics, and doubly special relativity, and calculated several magnitudes, including the Lamb shift and the tunnel current. A detailed study of the Schrödinger equation, with the harmonic oscillator problem as a toy model, was carried out in the work of Kempf and his collaborators Kempf . The one-dimensional box problems and of the free particle within the GUP approach by Nozari and Azizi Nozari , work in which they proposed an approximation to the associated Schrödinger equation considering a fourth-order expression, reporting the analytical solution. The exact analytical solutions of the hydrogen atom were obtained by Brau Brau and the analysis of the inverse square term was performed by Bouaziz and Bawin in Bouaziz . The dispersion states of the Woods-Saxon potential were investigated in Maghsoodi . In the relativistic realm, Hassanabadi et al. Rajabi proposed an approximate scheme to study the Dirac equation in the GUP formalism based on the resolution of a Schrödinger-like counterpart. Furthermore, the concept of GUP has been considered in the study of the thermodynamics of black holes in very recent articles Bargueno ; Buoninfante ; Lambiase . It is worth noting that Bishop and his collaborators, in a series of articles, have commented on a very deep conceptual point within the GUP formalism: in Singleton it is shown that, contrary to the common thinking, “different pairs of modified operators can lead to the same modified commutator and still give a different minimum length or even none at all”, and Bishop explains exhaustively that the results depend on how the modified operators are defined and that the resulting modified commutator is not the only important aspect.
On the other hand, the Coulomb potential is obviously the most physical interaction in common with gravity and quantum mechanics. It has been shown, through various approaches Kirilina ; Netto ; Verch , that the quantum correction of this potential will contain inverse quadratic, inverse cubic and inverse quartic terms. This detail is of particular interest to us, since what we are going to do in this work is analyze this type of corrected Coulomb potential that includes negative powers up to the fourth order, which we will call Coulomb(-4) potential, within the GUP formalism.
This work is organized as follows. The Schrödinger equation modified with the GUP formalism is reviewed compactly in Section 2. Section 3 investigates the solutions to the ordinary case (without GUP) using the Lie algebraic approach and Heun functions. In Section 4 both the ground state and the first excited state of the modified problem due to GUP are explicitly determined. The final section presents the conclusions and shows some of the possible lines of work that will be addressed in future work in this same area.
II The modified Schrödinger equation with a GUP and a quantum correction to the Coulomb interaction
We consider a GUP of the form Kempf
$\left[x_{G},p_{G}\right]=i\hbar\left(1+\frac{\beta}{2\hbar^{2}}\ p^{2}\right)\,,$ (1)
where the parameter $0\leq\beta\leq 1$ comes from the associated GUP, a change of coefficient is considered in the original form of the commutation relation for further convenience and the generalized $x$ operator is defined as
$x_{G}=x,$ (2)
with $x$ and $p$ being the ordinary position and momentum operators, respectively. In one spatial dimension, and neglecting some high order terms, the above GUP corresponds to the modified Schrödinger equation Maghsoodi
$\left(-\frac{d^{2}}{dx^{2}}+V_{e}(x)\right)\psi_{n}^{G}(x)=0,$ (3)
with an effective potential
$V_{e}(x):=\frac{2m}{\hbar^{2}}\left(V(x)-E_{n}^{G}\right)+\left(\frac{2m}{\hbar^{2}}\right)^{2}\beta\left(V(x)-E_{n}\right)^{2},$ (4)
in which $E_{n}$ denotes the ordinary case eigenvalues of the energy (with $\beta=0$) and $E_{n}^{G}$ the eigenenergies for $\beta\neq 0$. Although it is not explicitly indicated, obviously $E_{n}^{G}(\beta)$ are some functions of $\beta$ such that and $E_{n}^{G}(0)=E_{n}$. Note that, Eq. (3) is an approximation to the complete differential equation which appears, which is of order six Rajabi ; Nozari .
Now we are interested in analyzing a specific problem that could be of great physical interest. Based on Refs. Lambiase ; Kirilina ; Netto ; Verch , let us now consider the following quantum correction to the Coulomb interaction that includes negative powers up to fourth order, that we will call Coulomb(-4) potential:
$\frac{2m}{\hbar^{2}}V(x)=\frac{\alpha_{4}^{2}}{x^{4}}+\frac{\alpha_{3}}{x^{3}}+\frac{\alpha_{2}}{x^{2}}+\frac{\alpha_{1}}{x},\quad\alpha_{4}>0,\ \alpha_{1}<0,$ (5)
which corresponds to the effective potential
$V_{e}(x)=\frac{\gamma_{8}}{x^{8}}+\frac{\gamma_{7}}{x^{7}}+\frac{\gamma_{6}}{x^{6}}+\frac{\gamma_{5}}{x^{5}}+\frac{\gamma_{4}}{x^{4}}+\frac{\gamma_{3}}{x^{3}}+\frac{\gamma_{2}}{x^{2}}+\frac{\gamma_{1}}{x}+\gamma_{0}\,,$ (6)
in which the $\gamma_{i}$, $i=0,1,\dots 8$ are as follows
[…]
$\gamma_{1}$ $=\alpha_{1}\big{(}1-2\beta\epsilon_{n}\big{)},$ (7b)
$\gamma_{2}$ $=\alpha_{2}+\beta\big{(}\alpha_{1}^{2}-2\alpha_{2}\epsilon_{n}\big{)},$ (7c)
$\gamma_{3}$ $=\alpha_{3}+2\beta\left(\alpha_{1}\alpha_{2}-\alpha_{3}\epsilon_{n}\right),$ (7d)
$\gamma_{4}$ $=\alpha_{4}^{2}+\beta\big{(}2\alpha_{1}\alpha_{3}+\alpha_{2}^{2}-2\alpha_{4}^{2}\epsilon_{n}\big{)},$ (7e)
$\gamma_{5}$ $=2\beta(\alpha_{1}\alpha_{4}^{2}+\alpha_{2}\alpha_{3}),$ (7f)
$\gamma_{6}$ $=\beta\left(2\alpha_{2}\alpha_{4}^{2}+\alpha_{3}^{2}\right),$ (7g)
$\gamma_{7}$ $=2\beta\alpha_{3}\alpha_{4}^{2},$ (7h)
$\gamma_{8}$ $=\beta\alpha_{4}^{4},$ (7i)
where we have introduced the notation
$\epsilon_{n}=\frac{2m}{\hbar^{2}}\,E_{n},\qquad\epsilon_{n}^{G}=\frac{2m}{\hbar^{2}}\,E_{n}^{G}.$ (8)
It should be noted that, the problem here is considered in half-line where $x$ is positive and in the other direction an inpenetrable infinite wall is assumed. Since solutions of the model with GUP depend on the solutions of the associated ordinary case ($\beta=0$), let us first start with the analysis of this one.
III Schrödinger equation for a Coulomb(-4) potential
The ordinary case has already been studied via the ansatz method in Dong and also in AgZh2013 ; PaBar16 with some additional terms in the potential. Here, we will obtain the general solutions of the model using the Lie-algebraic method within the framework of quasi-exact solvability. Moreover, we will show that under appropriate transformation, the corresponding differential equation can be expressed in the form of the double-confluent Heun (DCH) equation. The Schrödinger equation with potential (5) appears in the form
$\left(-\frac{d^{2}}{dx^{2}}+\frac{\alpha_{4}^{2}}{x^{4}}+\frac{\alpha_{3}}{x^{3}}+\frac{\alpha_{2}}{x^{2}}+\frac{\alpha_{1}}{x}-\epsilon_{n}\right)\psi_{n}(x)=0.$ (9)
By inspecting the asymptotic behavior of the wave function $\psi_{n}(x)$, it follows that it is convenient to use the following ansatz Dong ; AgZh2013 ; Turbiner ; Turbiner88
$\psi_{n}(x)=x^{\delta}\,\exp\left[-\left(x\sqrt{-\epsilon_{n}}+\frac{\alpha_{4}}{x}\right)\right]\,\varphi_{n}(x),\qquad\delta=1+\frac{\alpha_{3}}{2\alpha_{4}}>0,$ (10)
which transforms Schrödinger equation (9) into the form
$\left\{-x^{2}\frac{d^{2}}{dx^{2}}-{2}\left(\alpha_{4}+\delta\,x-\sqrt{-\epsilon_{n}}\,x^{2}\right)\frac{d}{dx}+\left(\lambda_{1}\,x+\lambda_{2}\right)\right\}\varphi_{n}(x)=0,$ (11)
where
$\lambda_{1}=\alpha_{1}+\left(2+\frac{\alpha_{3}}{\alpha_{4}}\right)\sqrt{-\epsilon_{n}},\qquad\lambda_{2}=\alpha_{2}-\frac{\alpha_{3}^{2}}{4\alpha_{4}^{2}}-\frac{\alpha_{3}}{2\alpha_{4}}+2\alpha_{4}\sqrt{-\epsilon_{n}}.$ (12)
III.1 The Lie-algebraic approach
Following the standard idea of quasi-exact solvability PaBar16 ; Turbiner ; Turbiner88 ; Artemio94 ; BarPa18 , we find that if the constraint
$\lambda_{1}=-2n\sqrt{-\epsilon_{n}}$ (13)
holds, the equation (11) can then be expressed as a quasi-exactly solvable (QES) differential operator in the Lie-algebraic form AgZh2013 ; Turbiner ; Turbiner88
$H_{qes}\;\varphi_{n}(x)=0,$ (14)
with
$H_{qes}=-\mathcal{J}_{n}^{+}\mathcal{J}_{n}^{-}+2\sqrt{-\epsilon_{n}}\,\mathcal{J}_{n}^{+}-2\alpha_{4}\,\mathcal{J}_{n}^{-}-\Big{(}2\delta+n\Big{)}\mathcal{J}_{n}^{0}-\frac{n^{2}}{2}-n\delta+\lambda_{2}\,.$ (15)
Here,
$\mathcal{J}_{n}^{+}=x^{2}\,\frac{d}{dx}-n\,x,\qquad\mathcal{J}_{n}^{0}=x\,\frac{d}{dx}-\frac{n}{2},\ \qquad\mathcal{J}_{n}^{-}=\frac{d}{dx},$ (16)
are the generators of the $sl(2)$ Lie algebra satisfying the commutation relations PaBar16 ; Turbiner ; Artemio94 ; Turbiner88 ; Ronveaux
$[\mathcal{J}_{n}^{+},\mathcal{J}_{n}^{-}]=-2\mathcal{J}_{n}^{0},\qquad[\mathcal{J}_{n}^{\pm},\mathcal{J}_{n}^{0}]=\mp\mathcal{J}_{n}^{\pm}.$ (17)
The operators leave invariant the $(n+1)$-dimensional linear space of polynomials spanned by $\langle 1,x,x^{2},\dots,x^{n}\rangle$. In other words,
$\varphi_{n}(x)=\sum_{k=0}^{n}c_{k}\,x^{k},$ (18)
where the coefficients $c_{k}$ satisfy the three-term recursion relation
$c_{k+1}=\frac{\left(\lambda_{2}-2k\,\delta-k(k-1)\right)c_{k}-4\sqrt{-\epsilon_{n}}\,c_{k-1}}{2(k+1)\,\alpha_{4}},\qquad n=0,1,\dots,n,$ (19)
assuming $c_{-1}=0$ and $c_{n+1}=0$. Equivalently, the recursion relation (19) can be rewritten as a tridiagonal matrix equation the nontrivial solutions of which are determined by
$\left|\begin{array}[]{ccccc}\lambda_{2}&-2\alpha_{4}&&&\\[4.30554pt] -2n\sqrt{-\epsilon_{n}}&(\lambda_{2}-2\delta)&-4\alpha_{4}&&\\[4.30554pt] &-2(n-1)\sqrt{-\epsilon_{n}}&\ddots&\ddots&\\[4.30554pt] &&\ddots&\ddots&-2n\,\alpha_{4}\\[6.45831pt] &&&-2\sqrt{-\epsilon_{n}}&(n-n^{2}+\lambda_{2}-2n\,\delta)\end{array}\right|=0.$ (20)
Note that the condition (20) imposes severe restrictions on the potential parameters $\alpha_{i}\,$, $i=1,2,3,4$. On the other hand, from (12)–(13), the expression of energy in closed form can be obtained
$\epsilon_{n}=-\frac{\alpha_{1}^{2}\,\alpha_{4}^{2}}{(\alpha_{3}+2(n+1)\alpha_{4})^{2}},$ (21)
provided $\alpha_{1}<0$, which is concluded immediately after replacing $\lambda_{1}$ in (13) and taking into account that $\delta=1+\frac{\alpha_{3}}{2\alpha_{4}}>0$, as we have seen in (10).
Note that for the same reason, i.e. $\delta>0$, the expression in large square brackets is obviously non-zero. Therefore, for any given $n$, the general solutions of the energies and the corresponding wave functions, for the ordinary case, are given respectively by (21) and (10) together with (19) , as long as the restriction (20) is maintained on the parameters of the potential. Next, we represent the explicit ground state and first excited state solutions.
i. Ground state solution
From (21) and (10), the ground state energy and the associated wave function are given by
$\epsilon_{0}=-\frac{\alpha_{1}^{2}\,\alpha_{4}^{2}}{(\alpha_{3}+2\alpha_{4})^{2}},$ (22)
and
$\psi_{0}(x)=c_{0}\ x^{1+\alpha_{3}/(2\alpha_{4})}\ \exp\left[-\left(x\sqrt{-\epsilon_{0}}+\frac{\alpha_{4}}{x}\right)\right],$ (23)
where, from (20), the restriction on the parameters of the potential is determined by $\lambda_{2}=0$ in (12), or more explicitly, by
$\frac{\alpha_{3}}{\alpha_{4}}=1\pm\sqrt{1+4\alpha_{2}+8\alpha_{4}\sqrt{-\epsilon_{0}}}.$ (24)
ii. First excited state solution
For $n=1$, from (21) and (10), the energy and the corresponding wave function are given by
$\epsilon_{1}=-\frac{\alpha_{1}^{2}\,\alpha_{4}^{2}}{\left(\alpha_{3}+4\alpha_{4}\right)^{2}}\,,$ (25)
and
$\psi_{1}(x)=(c_{0}+c_{1}x)\ x^{1+\alpha_{3}/(2\alpha_{4})}\ \exp\left[-\left(x\sqrt{-\epsilon_{1}}+\frac{\alpha_{4}}{x}\right)\right],$ (26)
in which, from (19), we have
$c_{1}=\frac{\lambda_{2}}{2\alpha_{4}}\,c_{0}\,.$ (27)
In this case, from (20), the restriction on the parameters of the potential is given by
$\lambda_{2}(\lambda_{2}-2\delta)-4\alpha_{4}\,\sqrt{-\epsilon_{1}}=0,$ (28)
or explicitly, by
$64\alpha_{1}^{2}\alpha_{4}^{8}+16\alpha_{4}^{4}(\alpha_{3}+4\alpha_{4})(\alpha_{3}^{2}+4\alpha_{3}\alpha_{4}+8\alpha_{4}^{2}-4\alpha_{2}\alpha_{4}^{2})\alpha_{1}$
$\qquad+(\alpha_{3}+4\alpha_{4})^{2}(\alpha_{3}^{2}+2\alpha_{3}\alpha_{4}-4\alpha_{2}\alpha_{4}^{2})(\alpha_{3}^{2}+6\alpha_{3}\alpha_{4}+8\alpha_{4}^{2}-4\alpha_{2}\alpha_{4}^{2})=0.$ (29)
III.2 The ordinary case as a double-confluent Heun equation
Making the change of variable $y=2\sqrt{-\epsilon_{n}}\,x$, one finds that the differential equation (11) is transformed into a double-confluent Heun equation, namely Ronveaux
$\left\{y^{2}\frac{d^{2}}{dy^{2}}+\left(-y^{2}+\rho\,y+2\alpha_{4}\right)\frac{d}{dy}-\left(\omega\,y+\lambda_{2}\right)\right\}\varphi_{n}(y)=0,$ (30)
in which we denoted
$\rho=2+\frac{\alpha_{3}}{\alpha_{4}},\qquad\omega=1+\frac{\alpha_{1}}{2\sqrt{-\epsilon_{n}}}+\frac{\alpha_{3}}{\alpha_{4}},$ (31)
The regular solutions of (30) at origin, from Ronveaux ; Jaick , is given by
$\varphi:=\varphi_{n}(y\,;\rho,\alpha_{4},\omega,\lambda_{2})=\sum_{n}^{\infty}h_{n}\,x^{n},$ (32)
in which the coefficients $h_{n}$ satisfy the relations
$h_{n+2}=\frac{\left(\lambda_{2}-(n-\rho)(n+1)\right)h_{n+1}+\left(n-\omega\right)h_{n}}{-2\alpha_{4}}\qquad\text{and}\qquad h_{1}=\frac{\lambda_{2}}{2\alpha_{4}}\,h_{0}\,,$
assuming $h_{0}=1$. As a result, $\varphi$ can admit polynomial solution of degree $n$ if $(n-\omega)$ and $h_{n+1}$ vanish simultaneously. In this way, one can obtain the general solutions of the problem in terms of the associated Heun function. Needless to say, the results agree with those obtained using the Lie algebraic approach used in the previous section.
IV Solutions of the Coulomb(-4) model with a GUP
Having obtained the solutions of the Coulomb(-4) model in the previous section, let us now return to the modified Schrödinger equation in a GUP formalism, that is the equation (3). To ensure proper asymptotic behavior of the wave function $\psi_{n}^{G}(x)$, after inspecting the aforementioned differential equation (3) we propose the following ansatz
$\psi_{n}^{G}(x)=x^{f}\,e^{g(x)}\,\varphi_{n}^{G}(x),\qquad g(x)=-a\,x-\frac{b}{x}-\frac{c}{x^{2}}-\frac{d}{x^{3}},$ (33)
where the new function $\varphi_{n}^{G}(x)$ is a polynomial function and the parameters $a>0\,,b\,,c\,,d,\,f>0$ are still unknown, but must be such that the original function $\psi_{n}^{G}(x)$ is square integrable. Substituting (33) into (3) and solving the Riccati equation that results for $g(x)$, it can be seen that the previous parameters are determined by
$a$ $=\sqrt{\gamma_{0}}=\sqrt{\beta\epsilon_{n}^{2}-\epsilon_{n}^{G}}>0,$ (34a)
$b$ $=-\frac{\gamma_{7}^{2}-4\gamma_{6}\,\gamma_{8}}{8\,(\gamma_{8})^{3/2}}=\alpha_{2}\,\sqrt{\beta},$ (34b)
$c$ $=\frac{\gamma_{7}}{4\sqrt{\gamma_{8}}}=\frac{1}{2}\,\alpha_{3}\,\sqrt{\beta},$ (34c)
$d$ $=\frac{\sqrt{\gamma_{8}}}{3}=\frac{1}{3}\alpha_{4}^{2}\,\sqrt{\beta},$ (34d)
$f$ $=2+\frac{8\gamma_{5}\,\gamma_{8}^{2}-4\gamma_{6}\,\gamma_{7}\,\gamma_{8}+\gamma_{7}^{3}}{16\,(\gamma_{8})^{5/2}}=2+\alpha_{1}\sqrt{\beta}>0\,.$ (34e)
Then, the differential equation for $\varphi_{n}^{G}(x)$ simplifies to
$\left\{P_{4}(x)\frac{d^{2}}{dx^{2}}+Q_{4}(x)\frac{d}{dx}+W_{3}(x)\right\}\varphi_{n}^{GUP}(x)=0,$ (35)
where
$P_{4}(x)=$ $\ x^{4}\,,$ (36a)
$Q_{4}(x)=$ $\ 6d+4c\,x+2b\,x^{2}+2f\,x^{3}-2a\,x^{4}\,,$ (36b)
$W_{3}(x)=$ $\ (b^{2}-6ad+4cf-6c-\gamma_{4})+(2(bf-b-2ac)-\gamma_{3})x$
$\ +(f(f-1)-2ab-\gamma_{2})x^{2}-(\gamma_{1}+2af)x^{3},$ (36c)
the parameters $a\,,b\,,c\,,d,\,f$ and $\gamma_{i}\,$, $i=1,2,3,4$ are given respectively by (IV) and (II). To solve (35), we use the general Bethe ansatz method introduced in Appendix A in Ref. AgZh2013 . In this way, we look for polynomial solutions for $\varphi_{n}^{G}(x)$ of the form
$\varphi_{n}^{G}(x)=\!\left\{\begin{array}[]{cl}1,&\quad n=0,\\ \displaystyle\prod_{i=1}^{n}(x-x_{i}),&\quad n\in\mathbb{N},\end{array}\right.$ (37)
where $x_{i}$ are distinct roots to be determined. Applying then the Eqs. (A.6)-(A.10) from AgZh2013 , the ‘general solutions’ of (35) are given by this set of equations:
$2af+2an+\gamma_{1}=0,$ (38a)
$f(f-1)-2ab-\gamma_{2}-2a\sum_{i=1}^{n}x_{i}+2fn+(n-1)n=0,$ (38b)
$2bf-4ac-2b-\gamma_{3}-2a\sum_{i=1}^{n}x_{i}^{2}+2(f+n-1)\sum_{i=1}^{n}x_{i}+2n\,b=0,$ (38c)
$b^{2}-6ad+4cf-6c-\gamma_{4}-2a\sum_{i=1}^{n}x_{i}^{3}+2(f+n-1)\sum_{i=1}^{n}x_{i}^{2}+2\sum_{i<j}^{n}x_{i}\,x_{j}+2b\sum_{i=1}^{n}x_{i}+4n\,c=0,$ (38d)
where $x_{i}$ are the roots of the Bethe ansatz equations
$\sum_{j=1,\ j\neq i}^{n}\frac{1}{x_{i}-x_{j}}-\frac{a\,x_{i}^{4}-f\,x_{i}^{3}-b\,x_{i}^{2}-2c\,x_{i}-3d}{x_{i}^{4}}=0,\qquad i=1,2,\dots,n.$ (39)
The condition (38a) gives the energy relation: for a given $n$, the energy $\epsilon_{n}^{G}$ is determined by
$\alpha_{1}\left[1+2\sqrt{\beta(\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G})}-4\,\beta\,\epsilon_{n}\right]+2(n+2)\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}=0.$ (40)
Note that the term in square brackets is positive since $\epsilon_{n}<0$ times (21), again confirming that $\alpha_{1}$ must be negative for (40) to be satisfied. Through simple manipulations, (40) can be rewritten in closed form:
$\epsilon_{n}^{G}(\epsilon_{n};\alpha_{1},\beta)=-\frac{\alpha_{1}^{2}}{4}\left(\frac{2\beta\,\epsilon_{n}-1}{\alpha_{1}\,\sqrt{\beta}+(n+2)}\right)^{2}+\beta\,\epsilon_{n}^{2}\,.$ (41)
Let us keep in mind that the denominator of the fraction, $\alpha_{1}\,\sqrt{\beta}+(n+2)$, is always positive because $f>0$, as established in (IV), and remember that $\epsilon_{n}$ is given by (21). On the other hand, the equations (38b)-(38d) produce severe restrictions on the potential parameters, explicitly:
$\alpha_{2}$ $=\frac{2\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}\sum\limits^{n}_{i=1}x_{i}-\alpha_{1}\sqrt{\beta}(2n+3)-(n+1)(n+2)}{2\beta\,\epsilon_{n}-2\sqrt{\beta(\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G})}-1},$ (42a)
$\alpha_{3}$ $=2\,\frac{\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}\sum\limits^{n}_{i=1}x_{i}^{2}-\sqrt{\beta}\Bigl{(}(n+1)\alpha_{2}+\alpha_{1}\sum\limits^{n}_{i=1}x_{i}\Bigr{)}-(n+1)\sum\limits^{n}_{i=1}x_{i}}{2\beta\,\epsilon_{n}-2\sqrt{\beta(\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G})}-1},$ (42b)
$\alpha_{4}^{2}$ $=\frac{2\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}\sum\limits^{n}_{i=1}x_{i}^{3}-\sqrt{\beta}\Bigl{(}(2n+1)\alpha_{3}+2\alpha_{2}\sum\limits^{n}_{i=1}x_{i}+2\alpha_{1}\sum\limits^{n}_{i=1}x_{i}^{2}\Bigr{)}-2(n+1)\sum\limits^{n}_{i=1}x_{i}^{2}-2\sum\limits^{n}_{i<j}x_{i}\,x_{j}}{2\beta\,\epsilon_{n}-2\sqrt{\beta(\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G})}-1}\,.$ (42c)
The denominators of the fractions are again non-zero since $\epsilon_{n}<0$ by (21). The parameters $x_{i}$ are determined by the Bethe ansatz equations
$\sum_{j=1,\,j\neq i}^{n}\frac{1}{x_{i}-x_{j}}-\frac{x_{i}^{4}\sqrt{\beta\,\epsilon_{n}^{2}-\epsilon_{n}^{G}}-\sqrt{\beta}\left(\alpha_{4}+\alpha_{1}\,x_{i}^{3}+\alpha_{2}\,x_{i}^{2}+\alpha_{3}\,x_{i}\right)-2x_{i}^{3}}{x_{i}^{4}}=0,\quad i=1,2,\dots,n.$ (43)
In summary, for a given $n$, the ‘general solutions’ of the Schrödinger equation modified with a GUP are given by the equations (33) together with (37) and (41)-(43) As in the case of the ordinary model (without GUP), below we will look for explicit solutions for the ground state and the first excited state.
(i) Ground state
For $n=0$, from (41), the ground-state energy, $\epsilon_{0}^{G}$, is determined in closed form
$\epsilon_{0}^{G}(\epsilon_{0};\alpha_{1},\beta)=-\frac{\alpha_{1}^{2}}{4}\left(\frac{2\beta\,\epsilon_{0}-1}{\alpha_{1}\,\sqrt{\beta}+2}\right)^{2}+\beta\,\epsilon_{0}^{2}\,,$ (44)
where the ground-state energy of the ordinary case, $\epsilon_{0}$, was already obtained explicitly in (22). The explicit form of the associated wave function, from (33) together with (IV) and (37), is given by
$\psi_{0}^{G}(x)=C_{0}\ x^{2+\alpha_{1}\sqrt{\beta}}\ \exp\left[-x\,\sqrt{\beta\,\epsilon_{0}^{2}-\epsilon_{0}^{G}}-\frac{\sqrt{\beta}}{2}\left(\frac{2\alpha_{4}^{2}}{3x^{3}}+\frac{\alpha_{3}}{x^{2}}+\frac{2\alpha_{2}}{x}\right)\right],$ (45)
where $C_{0}$ is the normalization constant and, due to (IV), the potential parameters satisfy the constraints
$\alpha_{2}$ $=\frac{2+3\sqrt{\beta}\,\alpha_{1}}{2\sqrt{\beta\left(\beta\,\epsilon_{0}^{2}-\epsilon_{0}^{G}\right)}-2\beta\,\epsilon_{0}+1},$ (46a)
$\alpha_{3}$ $=\frac{2\sqrt{\beta}}{2\sqrt{\beta\left(\beta\,\epsilon_{0}^{2}-\epsilon_{0}^{G}\right)}-2\beta\,\epsilon_{0}+1}\ \alpha_{2},$ (46b)
$\alpha_{4}^{2}$ $=\frac{\sqrt{\beta}}{2\sqrt{\beta\left(\beta\,\epsilon_{0}^{2}-\epsilon_{0}^{G}\right)}-2\beta\,\epsilon_{0}+1}\ \alpha_{3}\,.$ (46c)
It is interesting to note that the Bethe ansatz (39) equations play no role in the ground state solution. In this simple case, substituting $\alpha_{2}$ and $\alpha_{3}$ into $\alpha_{4}$ in (IV), we find that $-\frac{2}{3\sqrt{\beta}}<\alpha_{1}<0$ so that $\alpha_{4}>0$ according to the assumption made in (5).
(ii) First excited state
For $n=1$, from (41), the ground state energy, $\epsilon_{1}^{G}$, is given by the closed form expression
$\epsilon_{1}^{G}(\epsilon_{1};\alpha_{1},\beta)=-\frac{\alpha_{1}^{2}}{4}\left(\frac{2\beta\,\epsilon_{1}-1}{\alpha_{1}\,\sqrt{\beta}+3}\right)^{2}+\beta\,\epsilon_{1}^{2}\,.$ (47)
provided $\alpha_{1}<0$. Recall again that $\epsilon_{1}$ was already obtained in (25). In the present case it is assumed that the differential equation (35) has solutions in the form (37), that is, $\varphi_{1}^{GUP}(x)=x-x_{1}\,$, and consequently, from (33)–(IV), the wave function is explicitly given by
$\psi_{1}^{G}(x)=C_{1}\ (x-x_{1})\ x^{2+\alpha_{1}\sqrt{\beta}}\ \exp\left[-x\,\sqrt{\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}}-\frac{\sqrt{\beta}}{2}\left(\frac{2\alpha_{4}^{2}}{3x^{3}}+\frac{\alpha_{3}}{x^{2}}+\frac{2\alpha_{2}}{x}\right)\right],$ (48)
where $C_{1}$ is the normalization constant. In addition, from (IV), we have the following constraints on the potential parameters
$\alpha_{2}$ $=\frac{2\,x_{1}\,\sqrt{\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}}-5\,\alpha_{1}\,\sqrt{\beta}-6}{-2\sqrt{\beta\left(\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}\right)}+2\beta\,\epsilon_{1}-1}\,,$ (49a)
$\alpha_{3}$ $=\frac{2\,x_{1}^{2}\,\sqrt{\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}}-2\sqrt{\beta}(2\alpha_{2}+\alpha_{1}\,x_{1})-4\,x_{1}}{-2\sqrt{\beta\left(\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}\right)}+2\beta\,\epsilon_{1}-1}\,,$ (49b)
$\alpha_{4}^{2}$ $=\frac{2\,x_{1}^{3}\,\sqrt{\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}}-\sqrt{\beta}\,\big{(}3\alpha_{3}+2(\alpha_{2}+\alpha_{1}\,x_{1})x_{1}\big{)}-4\,x_{1}^{2}}{-2\sqrt{\beta\left(\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}\right)}+2\beta\,\epsilon_{1}-1}\,.$ (49c)
The unknown parameter $x_{1}$ appearing in (48) and (IV) is determined from the Bethe ansatz equation (39), i.e.
$\sqrt{\beta\,\epsilon_{1}^{2}-\epsilon_{1}^{G}}\,x_{1}^{4}-\left(2+\alpha_{1}\sqrt{\beta}\right)\,x_{1}^{3}-\sqrt{\beta}\left(\alpha_{2}\,x_{1}^{2}+\alpha_{3}\,x_{1}+\alpha_{4}^{2}\right)=0\,.$ (50)
V Conclusions
Based on recent papers that propose a correction of the Coulomb potential by including quadratic, cubic and inverse quartic terms, we consider this modified potential in the framework of the modified Schrödinger equation with a generalized uncertainty principle. The resulting equation appears as a generalization of the Heun equation, that is, the biconfluent Heun equation that has the form $y^{\prime\prime}(x)+\left(A_{0}+\frac{A_{-1}}{x}+\frac{A_{-2}}{x^{2}}+\frac{A_{-3}}{x^{3}}+\frac{A_{-4}}{x^{4}}\right)y(x)=0$ Ronveaux . As far as we know, the general form we considered, could not be solved by other analytical techniques, including the common Lie-algebraic approach, which requires the form $P_{4}(x)\Phi^{\prime\prime}(x)+Q_{3}(x)\Phi^{\prime}(x)+W_{2}(x)\Phi(x)=0$ Turbiner ; Turbiner88 . Therefore, we try a Bethe-Ansatz to the equation and after some calculations we find the ground state and the first excited state. It is worth mentioning that despite its merits, this approach becomes longer for higher states as the parameters involved and their determination become more complicated, as some constraints are imposed on the parameters in order to provide the solutions. However, the truth is that this last limitation exists in most approaches, including supersymmetric quantum mechanics, power series solutions, integral transforms, etc. We are currently working on the generalization of these solutions to few-body systems, to problems in more dimensions and to other physical fields.
Acknowledgments
L.M.N. and S.Z. were supported by Spanish MCIN with funding from European Union Next Generation EU (PRTRC17.I1) and Consejeria de Educacion from JCyL through QCAYLE project, as well as grants PID2020-113406GB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T. The work of M.B. was supported by the Czech Science Foundation within the project 22-18739S.